Ben is twice as old as mike.three years ago,he was three times as old as mike. how old is each boy now?
Answer by nshah11(47)

B = Ben's Age
M = Mike's age
B = 2M
B - 3 = 3(M - 3)
2M - 3 = 3M - 9
M = 6
B = 12